# The Regular property of Invariant Rings over Regular Domains

### v2

@@ Abstract.
The main result of this paper is a generalization of the theorem of Chevalley-Shephard-Todd to the rings of invariants of pseudo-reflection groups over regular domains. More precisely, let $A$ be a regular domain and let $K$ be its field of fractions. Let $G\subseteq GL_{n}(A)$ be a finite group. Let $G$ act linearly on $A[X_{1},X_{2},\dots,X_{n}]$ (fixing $A$). Assume that $|G|$ is invertible in $A$. We prove that if $G\subseteq GL_{n}(K)$ is generated by psuedo-reflections then $(A[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular.
@@ 1. Introduction
$K[X_{1},X_{2},\dots,X_{n}]$ and let $G$ act linearly on the ring (fixing $K$). Assume that $|G|$ is invertible in $K$. We have $G\subseteq GL_{n}(K)$ is generated by psuedo-reflections if and only if $(K[X_{1},X_{2},\dots,X_{n}])^{G}$ is a polynomial ring over $K$.
@@
Let $A$ be a regular domain and let $K$ be its field of fractions. Consider a finite group $G\subseteq GL_{n}(A)$. Consider the ring $A[X_{1},X_{2},\dots,X_{n}]$ and let $G$ act linearly on the ring (fixing $A$). Assume that $|G|$ is invertible in $A$. If $G\subseteq GL_{n}(K)$ is generated by psuedo-reflections, then $(A[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular.
In Section 2, we prove Theorem 1.2 for the case when $A$ is a DVR as Theorem 2.2. This also leads to an alternate proof for the case of Dedekind domains considered by Mundelius in [2]. In Section 3, we prove Theorem 1.2 for the case when $A$ is a regular local ring as Theorem 3.2. Finally in Section 4, we complete the proof of Theorem 1.2.
@@
Let $(\mathcal{O},(\pi))$ be a DVR and $\mathcal{O}_{\pi}$ be the localization of $\mathcal{O}$ at the multiplicative set $\{\pi^{i}\}_{i}$ (which is the field of fractions of $\mathcal{O}$). Consider a finite group $G\subseteq GL_{n}(\mathcal{O})$. Consider the ring $\mathcal{O}[X_{1},X_{2},\dots,X_{n}]$ and let $G$ act linearly on the ring (fixing $\mathcal{O}$). Assume that $|G|$ is invertible in $\mathcal{O}$ i.e. $|G|$ is invertible in $\mathcal{O}/\pi\mathcal{O}$.
@@
Assume Hypothesis 2.1. Consider $G\subseteq GL_{n}(\mathcal{O}_{\pi})$ is generated by psuedo-reflections. Then $(\mathcal{O}[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular.
@@
Assume Hypothesis 2.1. Suppose that $G$ is generated by psuedo-reflections in $\mathcal{O}_{\pi}$. Then for $n\geq 2$ and any psuedo-reflection $\sigma\in G\subseteq GL_{n}(\mathcal{O}_{\pi})$, there exists a basis $\{w_{1},w_{2},\dots w_{n}\}$ of $\mathcal{O}^{n}$ with $\sigma(w_{i})=w_{i}$ for $1\leq i\leq n-1$ and $\sigma(w_{n})=\lambda w_{n}$ where $\lambda$ is an $m$-th root of unity where $m=ord(\sigma)$.
@@
We prove the result by induction on $n\geq 2$. Let $n=2$ and let $\{v_{1},v_{2}\}$ be a basis of $\mathcal{O_{\pi}}^{2}$ such that $\sigma(v_{1})=v_{1}$ and $\sigma(v_{2})=\lambda v_{2}$. We can assume $v_{1},v_{2}\in\mathcal{O}^{2}\backslash\pi\mathcal{O}^{2}$. Note $\lambda\neq 1$. Let $w_{1}=v_{1}$ and let $\{w_{1},w_{2}\}$ be a basis of $\mathcal{O}^{2}$. Let $v_{2}=\alpha w_{1}+\beta w_{2}$. Note that $\beta\neq 0$. Now $\sigma(v_{2})=\lambda v_{2}=\lambda\alpha w_{1}+\lambda\beta w_{2}$. Also we have $\sigma(v_{2})=\alpha\sigma(w_{1})+\beta\sigma(w_{2})=\alpha w_{1}+\beta\sigma(w_{2})$. Thus comparing both equations we have
(2.1) $\beta\sigma(w_{2})=(\lambda-1)\alpha w_{1}+\lambda\beta w_{2}$
Let $\lambda>$ be the subgroup of $\mathcal{O}^{\times}$ generated by $\lambda$. Since $ord(\lambda)=ord(\sigma)$, we have $|ord(\lambda)|\mid r$. As $|G|$ is invertible in $\mathcal{O}/\pi\mathcal{O}$, so $|<\lambda>|$ is also invertible in $\mathcal{O}/\pi\mathcal{O}$. By Proposition 2.3, the map $\lambda>\rightarrow(\mathcal{O}/\pi\mathcal{O})^{\times}$ is injective. As $\lambda\neq 1$, we have that the residue of $\lambda$ in $\mathcal{O}/\pi\mathcal{O}$ is $\bar{\lambda}\neq 1$. Thus $\lambda-1\not\in\pi\mathcal{O}$. So $\lambda-1\in\mathcal{O}^{\times}$.
We claim that $\beta\in\mathcal{O}^{\times}$. Assume on the contrary $\beta\in\pi\mathcal{O}$. Since $\alpha w_{1}+\beta w_{2}=v_{2}\not\in\pi\mathcal{O}^{2}$. Thus $\alpha\not\in\pi\mathcal{O}$. Going modulo $\pi$ in the Equation 2.1, we have $(\bar{\lambda}-1)\bar{\alpha}\bar{w_{1}}=0$ which is a contradiction. Thus we can write Equation 2.1 as $\sigma(w_{2})=\lambda w_{2}+\theta w_{1}$ for some $\theta\in\mathcal{O}$. Then letting $w^{\prime}_{2}=\frac{\theta}{\lambda-1}w_{1}+w_{2}$, we have that $\{w_{1},w^{\prime}_{2}\}$ is a basis and $\sigma(w^{\prime}_{2})=\lambda w^{\prime}_{2}$.
Let $n\geq 3$ and $\{v_{1},v_{2},\dots,v_{n}\}$ be a basis of $\mathcal{O}_{\pi}$ with $v_{i}\in\mathcal{O}^{n}\backslash\pi\mathcal{O}^{n}$ and $\sigma(v_{i})=v_{i}$ for $1\leq i\leq n-1$ and $\sigma(v_{n})=\lambda v_{n}$ and $\lambda\neq 1$. Let $E=\mathcal{O}v_{1}$ and $w_{1}=v_{1}$. We have the following exact sequence.
$0\rightarrow E\rightarrow W=\mathcal{O}^{n}\rightarrow\bar{W}=\mathcal{O}^{n-1}\rightarrow 0.$
Now $\sigma|_{\bar{W}}$ is also a psuedo-reflection. By induction we have a basis $\{\bar{w_{2}},\bar{w_{3}},\dots,\bar{w_{n}}\}$ of $\bar{W}$ with $\sigma(\bar{w_{i}})=\bar{w_{i}}$ for $2\leq i\leq n-1$ and $\sigma(w_{n})=\lambda\bar{w_{n}}$. Now $\{w_{1},w_{2},\dots,w_{n}\}$ is a basis of $W$ and $\sigma(w_{1})=w_{1}$ and $\sigma(w_{i})=w_{i}+a_{i}w_{1}$ for $2\leq i\leq n-1$ and $\sigma(w_{n})=\lambda w_{n}+a_{n}w_{1}$. We will show that $a_{2}=a_{3}=\dots=a_{n-1}=0$. Consider $2\leq i\leq n-1$. Now $\sigma(w_{i})=w_{i}+a_{i}w_{1}$. Thus $\sigma^{j}(w_{i})=w_{i}+ja_{i}w_{1}$. Now $\sigma^{m}=id$ where $m=ord(\sigma)=ord(\lambda)$. Thus $w_{i}=w_{i}+ma_{i}w_{1}$. Hence $ma_{i}w_{1}=0$. As $m$ is invertible in $\mathcal{O}$, we have $a_{i}=0$. Also by Proposition 2.3, $\lambda-1$ is invertible in $\mathcal{O}$. Thus $w_{1},w_{2},\dots,w_{n-1},w^{\prime}_{n}$ is the required basis where $w^{\prime}_{n}=\frac{a_{n}}{\lambda-1}w_{1}+w_{n}$ and $\sigma(w^{\prime}_{n})=\lambda w^{\prime}_{n}$. ∎
@@
Assume Hypothesis 2.1. Suppose that $G$ is generated by psuedo-reflections in $\mathcal{O}_{\pi}$. Then for $n\geq 2$, the image of $G$ in $GL_{n}(\mathcal{O}/\pi\mathcal{O})$ is also generated by psuedo-reflections.
@@
Thus $R^{G}/\pi R^{G}\cong(R/\pi R)^{G}$. Now $R/\pi R=(\mathcal{O}/\pi\mathcal{O})[X_{1},X_{2},\dots,X_{n}]$ and $\mathcal{O}/\pi\mathcal{O}$ is a field. Also by Corollary 2.4.1 we have that $G\subseteq GL_{n}(\mathcal{O}/\pi\mathcal{O})$ is generated by psuedo-reflections. Thus by Theorem 1.1, $(R/\pi R)^{G}$ is regular. Let $\mathfrak{M}$ be the maximal homogenous ideal of $R^{G}$. Since $\pi\not\in\mathfrak{M}^{2}$. Thus $R^{G}/\pi R^{G}$ being regular implies that $S=R^{G}$ is regular.∎
@@
Let $(A,\mathfrak{m})$ be a regular local ring and let $K$ be its field of fractions. Consider a finite group $G\subseteq GL_{n}(A)$. Consider the ring $A[X_{1},X_{2},\dots,X_{n}]$ and let $G$ act linearly on the ring (fixing $A$). Assume that $|G|$ is invertible in $A$.
@@
Assume Hypothesis 3.1. Consider $G\subseteq GL_{n}(K)$ is generated by psuedo-reflections. Then $(A[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular.
@@
Consider the case in Hypothesis 3.1. Let $d=\dim(A)$. Now Theorem 3.2 for the case $d=0$ (i.e. $A$ is a field) follows from Theorem 1.1. Thus we can assume $d\geq 1$. Let $\mathfrak{m}=(x_{1},\dots,x_{d})$. Set $x=x_{1}$ and let $\mathfrak{p}$ be the prime ideal $(x)$.
@@
As $A$ is a regular local ring we have that $A_{\mathfrak{p}}$ is a regular local ring of dimension $1$, i.e. a DVR which has $K$ as its field of fractions. Let $\kappa(\mathfrak{p})=A_{\mathfrak{p}}/\mathfrak{p}A_{\mathfrak{p}}$ which is the field of fractions of $A/\mathfrak{p}$. Now $G\subseteq GL_{n}(K)$ is generated by psuedo-reflections. Thus by Corollary 2.4.1, we have that the image of $G$ in $GL_{n}(\kappa(\mathfrak{p}))$ is generated by psuedo-reflections. Note that we have a commutative diagram of ring homomorphisms,
@@
We have that $\gamma|_{G}:G\rightarrow GL_{n}(A/\mathfrak{p})$ is injective by Lemma 3.4. Thus we have that $\mu\circ\gamma|_{G}$ is injective. Now $\gamma^{\prime}\circ\iota|_{G}=\mu\circ\gamma|_{G}$. So the image of $G$ in $GL_{n}(\kappa(\mathfrak{p}))$ is generated by psuedo-reflections. By induction we have $(A/\mathfrak{p}\ [X_{1},X_{2},\dots X_{n}])^{G}$ to be regular. Let $R=A[X_{1},X_{2},\dots,X_{n}]$ and $S=R^{G}$. Consider the short exact sequence
@@
Acknowledgements: We would like to thank Prof. Gregor Kemper for making us aware of the work of David Mundelius.

### v3

@@ Abstract.
The main result of this paper is a generalization of the theorem of Chevalley-Shephard-Todd to the rings of invariants of pseudo-reflection groups over regular domains. More precisely, let $A$ be a regular domain and let $K$ be its field of fractions. Let $G\subseteq GL_{n}(A)$ be a finite group. Let $G$ act linearly on $A[X_{1},X_{2},\dots,X_{n}]$ (fixing $A$). Assume that $|G|$ is invertible in $A$. We prove that $G\subseteq GL_{n}(K)$ is generated by pseudo-reflections if and only if $(A[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular.
@@ 1. Introduction
We begin by recalling the notion of pseudo-reflections.
Definition 1.1 (§7.1, [1]).
Let $K$ be a field. An element $\sigma\in GL_{n}(K)$ is called a pseudo-reflection if $ord(\sigma)$ is finite and $dim(ker(\sigma-1))=n-1$, i.e. whose fixed points have codimension one.
@@
$K[X_{1},X_{2},\dots,X_{n}]$ and let $G$ act linearly on the ring (fixing $K$). Assume that $|G|$ is invertible in $K$. We have $G\subseteq GL_{n}(K)$ is generated by pseudo-reflections if and only if $(K[X_{1},X_{2},\dots,X_{n}])^{G}$ is a polynomial ring over $K$.
@@
Let $A$ be a regular domain and let $K$ be its field of fractions. Consider a finite group $G\subseteq GL_{n}(A)$. Consider the ring $A[X_{1},X_{2},\dots,X_{n}]$ and let $G$ act linearly on the ring (fixing $A$). Assume that $|G|$ is invertible in $A$. We have $G\subseteq GL_{n}(K)$ is generated by pseudo-reflections if and only if $(A[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular.
One direction of the above theorem follows from Theorem 1.2. Here’s the proof.
Proof.
Suppose $(A[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular. Now $G$ acts linearly on the ring $A[X_{1},X_{2},\dots,X_{n}]$ (fixing $A$) and $|G|$ is invertible in $A$. As $K$ is the field of fractions of $A$, so $G$ also acts linearly on the ring $K[X_{1},X_{2},\dots,X_{n}]$ (fixing $K$) and $|G|$ is invertible in $K$. Now as $(K[X_{1},X_{2},\dots,X_{n}])^{G}$ is a localization of $(A[X_{1},X_{2},\dots,X_{n}])^{G}$, we have that $(K[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular. Therefore by Theorem 5.3.3 in [4] and by Theorem 19.2 in [2], $(K[X_{1},X_{2},\dots,X_{n}])^{G}$ is a polynomial ring over the field $K$. Hence by Theorem 1.2, $G\subseteq GL_{n}(K)$ is generated by pseudo-reflections . ∎
The subsequent sections of the paper are dedicated to proving the other direction. In Section 2, we prove Theorem 1.3 for the case when $A$ is a DVR as Theorem 2.2. This also leads to an alternate proof for the case of Dedekind domains considered by Mundelius in [3]. In Section 3, we prove Theorem 1.3 for the case when $A$ is a regular local ring as Theorem 3.2. Finally in Section 4, we complete the proof of Theorem 1.3.
@@
Assume that $(\mathcal{O},(\pi))$ is a DVR and $\mathcal{O}_{\pi}$ is the localization of $\mathcal{O}$ at the multiplicative set $\{\pi^{i}\}_{i}$ (which is the field of fractions of $\mathcal{O}$). Consider a finite group $G\subseteq GL_{n}(\mathcal{O})$. Consider the ring $\mathcal{O}[X_{1},X_{2},\dots,X_{n}]$ and assume that $G$ acts linearly on the ring (fixing $\mathcal{O}$). Assume that $|G|$ is invertible in $\mathcal{O}$ i.e. $|G|$ is invertible in $\mathcal{O}/\pi\mathcal{O}$.
@@
Assume Hypothesis 2.1. Consider $G\subseteq GL_{n}(\mathcal{O}_{\pi})$ is generated by pseudo-reflections. Then $(\mathcal{O}[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular.
@@
Assume Hypothesis 2.1. Suppose that $G$ is generated by pseudo-reflections in
$GL_{n}(\mathcal{O}_{\pi})$. Then for $n\geq 2$ and any pseudo-reflection $\sigma\in G\subseteq GL_{n}(\mathcal{O}_{\pi})$, there exists a basis $\{w_{1},w_{2},\dots w_{n}\}$ of $\mathcal{O}^{n}$ with $\sigma(w_{i})=w_{i}$ for $1\leq i\leq n-1$ and $\sigma(w_{n})=\lambda w_{n}$ where $\lambda$ is an $m$-th root of unity where $m=ord(\sigma)$.
@@
Let $\{v_{1},\ldots v_{n}\}$ be a basis of $(\mathcal{O}_{\pi})^{n}$ such that $\sigma(v_{i})=v_{i}$ for $1\leq i\leq n-1$ and $\sigma(v_{n})=\lambda v_{n}$.
Let $\lambda>$ be the subgroup of $\mathcal{O}^{\times}$ generated by $\lambda$. Since $ord(\lambda)=ord(\sigma)$, we have $|ord(\lambda)|\mid|G|$. As $|G|$ is invertible in $\mathcal{O}/\pi\mathcal{O}$, so $|<\lambda>|$ is also invertible in $\mathcal{O}/\pi\mathcal{O}$. By Proposition 2.3, the map $\lambda>\rightarrow(\mathcal{O}/\pi\mathcal{O})^{\times}$ is injective. As $\lambda\neq 1$, we have that the residue of $\lambda$ in $\mathcal{O}/\pi\mathcal{O}$ is $\bar{\lambda}\neq 1$. Thus $\lambda-1\not\in\pi\mathcal{O}$. So $\lambda-1\in\mathcal{O}^{\times}$.
Consider the endomorphism $\epsilon=(\lambda-1)^{-1}(\sigma-id)\colon\mathcal{O}^{n}\rightarrow\mathcal{O}^{n}$. By considering $\epsilon_{\pi}$ it follows that $\epsilon$ is an idempotent. So $\mathcal{O}^{n}=\ker\epsilon\oplus\ker(1-\epsilon)$. We note that $\operatorname{rank}(\ker(\epsilon))=n-1$ and $\operatorname{rank}(\ker(1-\epsilon))=1$. On $\ker(\epsilon)$ the map $\sigma$ acts as identity while on $\ker(1-\epsilon)$ the map $\sigma$ acts as multiplication by $\lambda$. The result follows. ∎
@@
Assume Hypothesis 2.1. Suppose that $G$ is generated by pseudo-reflections in $GL_{n}(\mathcal{O}_{\pi})$. Then for $n\geq 2$, the image of $G$ in $GL_{n}(\mathcal{O}/\pi\mathcal{O})$ is also generated by pseudo-reflections.
@@
Thus $R^{G}/\pi R^{G}\cong(R/\pi R)^{G}$. Now $R/\pi R=(\mathcal{O}/\pi\mathcal{O})[X_{1},X_{2},\dots,X_{n}]$ and $\mathcal{O}/\pi\mathcal{O}$ is a field. Also by Corollary 2.4.1 we have that $G\subseteq GL_{n}(\mathcal{O}/\pi\mathcal{O})$ is generated by pseudo-reflections. Thus by Theorem 1.2, $(R/\pi R)^{G}$ is regular. Let $\mathfrak{M}$ be the maximal homogenous ideal of $R^{G}$. Since $\pi\not\in\mathfrak{M}^{2}$. Thus $R^{G}/\pi R^{G}$ being regular implies that $S=R^{G}$ is regular.∎
@@
Assume that $(A,\mathfrak{m})$ is a regular local ring and $K$ is its field of fractions. Consider a finite group $G\subseteq GL_{n}(A)$. Consider the ring $A[X_{1},X_{2},\dots,X_{n}]$ and assume that $G$ acts linearly on the ring (fixing $A$). Assume that $|G|$ is invertible in $A$.
@@
Assume Hypothesis 3.1. Consider $G\subseteq GL_{n}(K)$ is generated by pseudo-reflections. Then $(A[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular.
@@
Consider the assumptions in Hypothesis 3.1. Let $d=\dim(A)$. Now Theorem 3.2 for the case $d=0$ (i.e. $A$ is a field) follows from Theorem 1.2. Thus we can assume $d\geq 1$. Let $\mathfrak{m}=(x_{1},\dots,x_{d})$. Set $x=x_{1}$ and let $\mathfrak{p}$ be the prime ideal $(x)$.
@@
As $A$ is a regular local ring we have that $A_{\mathfrak{p}}$ is a regular local ring of dimension $1$, i.e. a DVR which has $K$ as its field of fractions. Let $\kappa(\mathfrak{p})=A_{\mathfrak{p}}/\mathfrak{p}A_{\mathfrak{p}}$ which is the field of fractions of $A/\mathfrak{p}$. Now $G\subseteq GL_{n}(K)$ is generated by pseudo-reflections. Thus by Corollary 2.4.1, we have that the image of $G$ in $GL_{n}(\kappa(\mathfrak{p}))$ is generated by pseudo-reflections. Note that we have a commutative diagram of ring homomorphisms,
@@
We have that $\gamma|_{G}:G\rightarrow GL_{n}(A/\mathfrak{p})$ is injective by Lemma 3.4. Thus we have that $\mu\circ\gamma|_{G}$ is injective. Now $\gamma^{\prime}\circ\iota|_{G}=\mu\circ\gamma|_{G}$. So the image of $G$ in $GL_{n}(\kappa(\mathfrak{p}))$ is generated by pseudo-reflections. By induction we have $(A/\mathfrak{p}\ [X_{1},X_{2},\dots X_{n}])^{G}$ is regular. Let $R=A[X_{1},X_{2},\dots,X_{n}]$ and $S=R^{G}$. Consider the short exact sequence
@@
5. Conflict of interest statement
The authors have no conflict of interest to declare that are relevant to this article.
6. Data availability statement
The manuscript has no associated data.
Acknowledgements: We would like to thank Prof. Gregor Kemper for making us aware of the work of David Mundelius. We also thank the two referees for many pertinent comments.